OpenStudy (anonymous):

√(X-8 ) + √X = -5.. WHAT IS X?

OpenStudy (anonymous):

Hi

hartnn (hartnn):

the question is \[\sqrt{x-8}+\sqrt x=-5\] right ? First, you bring one square root on other side, so, \[\sqrt{x-8}=-\sqrt x-5\] now, square both sides, and tell me what you get ?

OpenStudy (anonymous):

x - 8 = x - 25 ???

OpenStudy (anonymous):

Is this the answer? 17 = 0

hartnn (hartnn):

ummm. nopes, actually \[(a-b)^2=a^2-2ab+b^2\] so, squaring both sides of \(\sqrt{x-8}=-\sqrt x-5\) will give you \(x-8=x-10\sqrt{x}+25\) because \((\sqrt{variable})^2=variable.\) now , try to isolate \(\sqrt x\) from that equation., what you get ?

OpenStudy (anonymous):

Thank you for your time. I have already gotten the answer from someone else.

hartnn (hartnn):

ohh....okk. welcome ^_^